OpenStudy (anonymous):

Calculus help please! Fist time taking this class The number of flowers planted, t hours after starting at 8, is given N(t)=18t_6t^2-1.3t^3 A) how many flowers had been planted by 11am? B At what rate is the number of flowers planted changing at 11am?

OpenStudy (anonymous):

I am assuming 8am. 11 is 3 hours after 8 so we want to let t = 3. For A you substitute 3 into the original equation The B you substitute 3 into the derivative (or find a limit) or use any other method for finding instantaneous ROC.

OpenStudy (anonymous):

Does that help you? Or can you not find the instantaneous ROC

OpenStudy (anonymous):

yes thank you! if i need more help is it okay if i message you?

OpenStudy (anonymous):

If i'm still online then sure

OpenStudy (anonymous):

Can i get some help with part b?

OpenStudy (anonymous):

@Riddellikins

OpenStudy (anonymous):

What part do you need help with? Or where are you up to?

OpenStudy (anonymous):

deriving the equation

OpenStudy (anonymous):

for part b

OpenStudy (anonymous):

N(t)=18t_6t^2-1.3t^3

OpenStudy (anonymous):

Equation and draw buttons are not working for me for some reason. I will do it now :P

OpenStudy (anonymous):

Is that underscore between 18t and 6t^2 meant to be a subtract?

OpenStudy (anonymous):

sorry i made a mistake there the equation is 18t+6t^2-1.3t^3

OpenStudy (anonymous):

Okay, I can do it now :)

OpenStudy (anonymous):

is the derivative t=4.35897+3.07692 t?

OpenStudy (anonymous):

N(t) = 18t+6t^2-1.3t^3 I derive using differentiation (I think) To do this you multiply the exponent (of the variable, t in this case) by the coefficient for each term and then you decrease the exponent by 1. N'(t) = 18x1 t^(1-1) + 6x2 t^(2-1) - 1.3x3 t^(3-1) N'(t) = 18t^0 + 12t - 3.9t^2 N'(t) = 18 + 12t - 3.9t^2

OpenStudy (anonymous):

If you need help deriving then wolfram alpha is a great tool to check your answers: http://www.wolframalpha.com/input/?i=derivative+18t%2B6t%5E2-1.3t%5E3

OpenStudy (anonymous):

Omg your the best thank you for all your help!